# Measuring AI Ability to Complete Long Tasks

### v2

@@ C.1 Ways in which models have improved
    For example, on the interpret_data task, GPT-4o fails in this run due to an incorrect interpretation of the request to find “the tallest building in the world in 1972” as finding “the tallest building in the world built in 1972”, as opposed to the common sense interpretation that the task is to find the tallest building in the world in 1972. For example, it writes code like:
@@
    In contrast, o1-preview correctly interprets the request as finding the tallest building in the world built in or before 1972 (e.g. see this run).
C.2 What do they still do poorly?
Despite improvements, even current frontier models still struggle in substantial ways. In general, our observerations indicate that agents tend to do worse on tasks that are qualitatively “messier”. We investigate this claim further in Section 6.2.
@@ C.2 What do they still do poorly?
In a capture the flag task, agents have to open a file with unknown encoding in python, and often fail on the first try because they specify an incorrect encoding. Even the best models respond by trying out different encodings in the python script as in this run, which is inefficient and wastes tokens. We never saw an agent running the bash file command that would identify the necessary encoding directly.
Appendix D More ablations and robustness checks
@@
    Alternative curve-fits (D.1)
@@
    Different weightings of task families: 1family $\frac{1}{\sqrt{\text{family size}}}$ vs uniform (included in Figure 12); also 1family $\frac{1}{\text{family size}}$
@@
D.1 Alternative curve fits
@@
In our main result, we fit an exponential curve (linear with a log y axis) because the fit is very good ($R^{2}\geq 0.96$, depending on the exact data and methodology). Linear and hyperbolic curves have poor fits (Figure 17). Because there are only 11 frontier models in the time span we studied, and the exponential fit has such high $R^{2}$ with only two parameters, we think applying fits with more parameters would be more likely to overfit than to give accurate predictions. In particular, we considered:
@@
  • Sophisticated stochastic models that asymptote to infinity often have large uncertainties; see Roodman [2020], which applied a superexponential diffusion model to economic data. As we discuss in Section 7.2, large uncertainty may be appropriate when constructing a prediction interval for a long-term forecast, but quantitatively modeling the potential impact of the factors we mentioned is out of scope of this paper, so we prefer to discuss these qualitatively.
@@
D.2 2024–2025 Horizon growth trend
@@ D.3 SWE-bench Verified
SWE-bench Verified contains time estimates created by contractors that split tasks into four buckets: $<15$ min fix, 15 minutes–1 hour, 1 hour–4 hours, or $>4$ hours. These estimates were based on the expected time it would take an “engineer who has had a few hours to familiarize themselves with the codebase” to solve the issue.
We verified annotator time buckets by running seven baselines across 6 SWE-bench Verified tasks. We baselined four tasks in the “$<15$ min fix” bucket, and found that they took our baseliners 8, 26, 67, and 84 minutes respectively. We baselined two tasks in the “1 hour–4 hours” bucket, and both baseliners took between 2–3 hours.
@@
$>4$ hours 480.0 min
Table 8: We convert the SWE-bench Verified time annotations into task estimates, by taking the geometric mean of the time annotation. We caution that this likely underestimates the time each issue takes to resolve by a human baseliner without context—notably, we observe that the geometric mean of baseliner time for four randomly sampled tasks in the “$<15$ minute fix” bucket is 32.9 minutes.
@@
As shown in Table 8, we found that annotator time estimates are likely more inaccurate on the “$<$15 min fix” time bucket, as in practice baseliners often take much longer than this to complete tasks. This results in easier tasks’ difficulty being underestimated, meaning that the SWE-bench Verified time horizon estimates likely underestimate true model performance for less capable models, which in turn leads to a steeper exponential trend.
@@
D.4 Messiness factors
Messiness factor labels were collected from 2 METR contractors and one researcher.
@@ D.4 Messiness factors
Not purely automatic scoring Does this task NOT have an entirely automatic scoring function? Specifically, for tasks implemented in the METR task standard: 1. There are no branches in which the scoring function returns None AND 2. The task provides manual scoring instructions in those cases AND 3. There are no clauses about obeying rules that are not actually checked in the automatic scoring. (e.g A family which had the instruction ”you may not attempt to reverse engineer X’” but this is not checked in the automatic scoring function)
@@
D.5 Randomizing file names
@@ D.5.2 Results
D.6 Correlation in performance between agents
In Figure 22, we report the correlation matrix of per-task success rates between each pair of models. In Figure 23, we report the correlation matrix for the excess success rate $({S_{observed}-S_{predicted}})/{S_{predicted}}$ (where $S_{predicted}$ is the success rate predicted from the model’s time horizon). While the correlations for excess success rate are lower than the correlation of raw task success rate (0.40 instead of 0.73), the fact that it is not zero suggests additional factors that explain model success rates across tasks which are consistent across models.
Figure 22: Correlation matrix of observed success rates across all models and tasks. Mean correlation: 0.73
Figure 23: Correlation matrix of excess success rates (defined by $\frac{S_{observed}-S_{predicted}}{S_{predicted}}$) across all models and tasks. Mean correlation: 0.40
@@ D.6 Correlation in performance between agents
Appendix E Code
Code and data to reproduce the core figures in this paper can be found at https://github.com/METR/eval-analysis-public.
Appendix F Author contributions

### v3

@@ D.2.1 What do newer models do better?
    For example, on the interpret_data task, GPT-4o fails in one run due to an incorrect interpretation of the request to find “the tallest building in the world in 1972” as finding “the tallest building in the world built in 1972”, as opposed to the common sense interpretation that the task is to find the tallest building in the world in 1972. For example, it writes code like:
@@
    In contrast, o1-preview correctly interprets the request as finding the tallest building in the world built in or before 1972.
D.2.2 What do they still do poorly?
Despite improvements, even current frontier models still struggle in substantial ways. In general, our observerations indicate that agents tend to do worse on tasks that are qualitatively “messier”. We investigate this claim further in Section F.2.
@@ D.2.2 What do they still do poorly?
In a capture the flag task, agents have to open a file with unknown encoding in python, and often fail on the first try because they specify an incorrect encoding. Even the best models respond by trying out different encodings in the python script, which is inefficient and wastes tokens. We never saw an agent running the bash file command that would identify the necessary encoding directly.
Appendix E Further discussion
E.1 Constant factors versus slope changes
Although extrapolation is always imperfect, forecasting AI time horizons far into the future is much more sensitive to changes in doubling rate than to constant factors in time horizon. For example, a naive extrapolation based on o1’s time horizon of 39 minutes and a past doubling time of 218 days (3.2x/year) predicts that AIs will reach a 1-month 50% time horizon (roughly 8 doublings over o1) about 4.8 years after the release date of o1. A 2x increase in doubling time would delay the 1 month point by a further 4.8 years, but a 2x constant factor decrease in time horizon would only delay it by 0.6 years.
E.2 Important factors that could change time horizon slope
Agency training
Compute scaling
Between the release of GPT-2 and today, the compute used to train the most impressive frontier language models has increased by at least a factor of 10,000x Epoch AI [2024], with training compute usage doubling every 6–10 months Sevilla et al. [2022]. More recently, models like o1 and o3 have began to use more compute at inference time. It is unclear whether there is sufficient capacity to expand either training or inference compute by many more orders of magnitude in the next 5 years. However, algorithmic improvements, which have historically decreased the compute requirements for a fixed performance level [Erdil and Besiroglu, 2023] [Ho et al., 2024], can substitute for compute limitations. We think that limits to compute scaling will slow the growth of AI agent time horizons somewhat, but be partially compensated by more investment into algorithmic improvement.
Automation of AI R&D
AGI will have “infinite” time horizon
An infinite time horizon does not mean an arbitrarily capable AI, merely the ability to complete tasks that take humans an arbitrarily long length of time. If an artificial general intelligence (AGI) is capable of completing all tasks expert humans can with a success rate of at least X%, its X% time horizon will necessarily be infinite. Therefore, if such systems are ever developed, the long-term trend in time horizon will be faster than exponential, with an asymptote at the date of AGI deployment.
E.3 Interpreting time horizon
Context and skill effects
At real companies, junior software engineer hires often take weeks of onboarding to begin contributing economic value. The human baseliners that determine the length of most tasks have much less context than average employees, potentially increasing measured task length. Our tasks are designed to require minimal context, which somewhat mitigates this problem; our internal PRs (Section C.2) were not designed this way, and so baseliners took many times longer than employees. However, highly skilled baseliners can also complete tasks far faster than average employees. Our expert baseliners are likely much more skilled than the average software engineer, potentially decreasing our measured task length.
Task distribution effects
Measuring extreme time horizons
Human time horizon measurements
In theory, one could also measure the time horizon of a human or population of humans. However, there are both theoretical and practical difficulties to doing so. We discuss this more in Appendix C.1.4.
E.4 Limitations and future work
More models with better elicitation
In general, we find that properly eliciting models can make a very large difference in their performance. [^17] This is a common obervation; see e.g. the improvements to software development capabilities from AIDE [Jiang et al., 2025] or recent work on KernelBench: [METR, 2025] We have put a limited amount of effort into eliciting models to get good performance on our tasks, so while our results are a reasonable lower bound, some models may have somewhat greater capabilities than we demonstrate. The most work was done to elicit o1 and the original Claude 3.5 Sonnet, each of which had around 2-3 engineer weeks of iterative development. All other models use the same scaffolding with at most minor changes. Future work could replicate our results with more effort spent on eliciting the full capabilities of frontier models.
More rigorous human baselining
Our per-task human time estimates are likely noisy due to relatively small sample size, and potentially also systematically skewed in various ways. Most notably, we select only successful completions of a task, and encourage baseliners to give up on tasks they may not complete in a reasonable amount of time. Our baseliners’ skills also vary significantly, and a wide variety of skills are relevant to these tasks. Though we attempt to match baseliners with appropriate tasks, this process is unlikely to be perfect. From manual reviews of baseline attempts, we also observe that humans sometimes simply give up even when the task seems within their capabilities, and it is unclear what selection effects are produced on the distribution of success times as a result. See See Appendix C.1.4 for further discussion of how bias in human baseline times could affect our results. Future work could replicate our results with more rigorous human baseliner selection or explore how sensitive the results are to methodological choices around human baselining.
More natural, varied tasks
There are reasons to believe that our task distribution is systematically different from the distribution of economically valuable work (and perhaps systematically different than the distribution of risk-model relevant tasks). We explored some of these reasons in Sections 4, but there remain many differences that we did not explore. For example, the modality of interaction in these tasks is also relatively narrow—for example, no tasks require the use of a mouse. No tasks require cooperating or competing with humans or other agents, [^18] See Xu et al. [2024] for an example of a recent benchmark that requires multi-agent interaction in a relatively realistic setting. while real software engineering or ML research involves communicating and coordinating with managers and other engineers or researchers. Many real-world tasks require very high reliability, and these are underrepresented in our dataset due to the difficulty of measuring models on these tasks. Most importantly, the tasks we study are heavily skewed toward software engineering and ML research. Future work could explore how the capabilities of AI agents are progressing in other domains.
More use of inference compute
Our scaffolds made relatively limited use of inference-time compute. When assuming that the human expert is compensated at \$143.61/hour (average L4 Engineer salary at Google divided by 2,000 hours), more than 80% of successful runs cost less than 10% of what it would cost for a human to perform the same task. (Figure 14). This implies that if inference-time computation could be used to improve performance, there is substantial room to do so while still remaining economically competitive with human experts. Previous research has found that techniques such as best-of-k can substantially improve performance on a subset of these tasks,Wijk et al. [2024] and better use of inference-time compute may lead to substantially different scores.
Data analysis
Appendix F External validity details
F.1 Retrodiction from 2023–2025 data
As part of exploratory work for this paper, we measured the time horizon of 9 frontier and near-frontier models released in 2023 and 2024, using only our HCAST and RE-Bench suites. This trend (with Claude 3.7 Sonnet and o3 added) is shown in Figure 10: time horizon doubles about every six months. Since we only had two 2023 models (GPT-4 0314 and GPT-4 1106) and a small data range (release date spanning 2 years and time horizon spanning 5 doublings), error bars were very wide. In addition, restricting our data further to 2024-only models produced a different trend with time horizon doubling about every three months, so any extrapolation into the future would not be robust.
To address these issues, we collected more data to extend the trendline into the past, developing the Software Atomic Actions (SWAA) suite to decrease the minimum human time of the task suite from 1 minute to under 2 seconds, and enabling us to measure GPT-2, davinci-002 (GPT-3) and GPT-3.5-turbo-instruct on the combined suite.
The 2023–2025 trend may be faster than the longer-term trend (Figure 11). The measured doubling time over the whole 6 year period 2019–2025 inclusive was 207 days, which is less steep than the 172-day trend based on data from non-SWAA tasks and 2023–2025 models. However, this difference is still within margin of error.
F.2 Messiness factors
Real-world intellectual labor often involves messy details that benchmarks usually don’t include, such as being under-specified or poorly scoped, having unclear feedback loops or success criteria, or requiring coordination between multiple streams of work in real-time. We generally observed that agents struggle more on tasks that have these “messy” details (Section 3.3). A natural question is therefore whether agents showed similar rates of improvement on “less messy” and “more messy” tasks.
We rated HCAST and RE-Bench tasks on 16 properties that we expected to be 1) representative of how real world tasks might be systematically harder than tasks we study and 2) relevant to AI agent performance. Some example factors include whether the task involved a novel situation, was constrained by a finite resource, involved real-time coordination, or was sourced from a real-world context. We labeled RE-bench and HCAST tasks on the presence or absence of these 16 messiness factors, then summed these to obtain a “messiness score” ranging from 0 to 16. Factor definitions can be found in Appendix H.6.
However, trends in AI agent performance over time are similar for lower and higher messiness subsets of the task suite. For example, on sub hour tasks, success rates increased by 40 percentage points between Jan. 2023 and May 2025 in both high and low messiness splits (Figure 12). In particular, we find no evidence of either much slower performance trends, or a plateau, specific to our higher messiness subset.
Appendix G Miscellaneous tables and figures
The following are several figures not included in the main body, from external validity and other sections.
Table 7: Average Success Rate by Model and Task Source
Task Source HCAST RE-Bench SWAA
Claude 3 Opus 0.25 0.00 0.98
Claude 3.5 Sonnet (New) 0.46 0.11 0.99
Claude 3.5 Sonnet (Old) 0.38 0.05 1.00
Claude 3.7 Sonnet 0.58 0.04 1.00
GPT-2 - - 0.40
GPT-4 0314 0.23 0.00 0.98
GPT-4 1106 0.30 0.02 0.97
GPT-4 Turbo 0.23 0.00 1.00
GPT-4o 0.30 0.00 0.98
davinci-002 (GPT-3) 0.00 0.00 0.65
gpt-3.5-turbo-instruct 0.01 0.00 0.95
o1 0.54 0.07 1.00
o1-preview 0.44 0.02 1.00
o3 0.65 0.40 -
o4-mini 0.61 0.27 -
Figure 12: Performance trends over time for HCAST and RE-Bench tasks by length and messiness (Section F.2). The data spans only 2023–2025 as pre-2023 models score 0 on non-SWAA tasks. Whilst our messier tasks have lower average success rates, trends in model performance improvements are not obviously slower on the high messiness split.
Figure 13: We plot the excess success rate (the observed empirical task success rate, minus success rate we would predict using the task’s length, see Section 3.1) against messiness score for each task. As discussed in Section F.2, there is a negative relationship between excess success rates and messiness.
Appendix H More ablations and robustness checks
@@
    Alternative curve-fits (H.1)
@@
    Different weightings of task families: $\frac{1}{\sqrt{\text{family size}}}$ vs uniform (included in Figure 6); also $\frac{1}{\text{family size}}$
@@
H.1 Alternative curve fits
@@
In our main result, we fit an exponential curve (linear with a log y axis) because the fit is very good ($R^{2}\geq 0.96$, depending on the exact data and methodology). Linear and hyperbolic curves have poor fits (Figure 15). Because there are only 12 frontier models in the time span we studied, and the exponential fit has such high $R^{2}$ with only two parameters, we think applying fits with more parameters would be more likely to overfit than to give accurate predictions. In particular, we considered:
@@
  • Sophisticated stochastic models that asymptote to infinity often have large uncertainties; see Roodman [2020], which applied a superexponential diffusion model to economic data. As we discuss in Section 5, large uncertainty may be appropriate when constructing a prediction interval for a long-term forecast, but quantitatively modeling the potential impact of the factors we mentioned is out of scope of this paper, so we prefer to discuss these qualitatively.
@@
H.2 Horizons at 80% success rate
In Figure 17, we show the trend for the 80% time horizon for models over time. The doubling time is similar to the 50% plot, but horizons are substantially lower.
H.3 Success rate versus inference cost
In Figure 18, we plot the success rate of models across our tasks as a function of token costs. All of the models studied perform better with larger token budgets, but the majority of models show clear evidence of plateaus far below the maximum number of tokens allocated to each model.
Figure 18: Success rate by cost. Models were given sufficiently high token limits to reach a plateau in success rate. Cost information is not included for o3 and o4-mini.
H.4 2024–2025 Horizon growth trend
@@ H.5 SWE-bench Verified
SWE-bench Verified contains time estimates created by contractors that split tasks into four buckets: <15<15 min fix, 15 minutes–1 hour, 1 hour–4 hours, or >4>4 hours. These estimates were based on the expected time it would take an “engineer who has had a few hours to familiarize themselves with the codebase” to solve the issue.
We verified annotator time buckets by running seven baselines across 6 SWE-bench Verified tasks. We baselined four tasks in the “<15<15 min fix” bucket, and found that they took our baseliners 8, 26, 67, and 84 minutes respectively. We baselined two tasks in the “1 hour–4 hours” bucket, and both baseliners took between 2–3 hours.
@@
>4>4 hours 480.0 min
Table 8: We convert the SWE-bench Verified time annotations into task estimates, by taking the geometric mean of the time annotation. We caution that this likely underestimates the time each issue takes to resolve by a human baseliner without context—notably, we observe that the geometric mean of baseliner time for four randomly sampled tasks in the “<15<15 minute fix” bucket is 32.9 minutes.
@@
As shown in Table 8, we found that annotator time estimates are likely more inaccurate on the “<<15 min fix” time bucket, as in practice baseliners often take much longer than this to complete tasks. This results in easier tasks’ difficulty being underestimated, meaning that the SWE-bench Verified time horizon estimates likely underestimate true model performance for less capable models, which in turn leads to a steeper exponential trend.
@@
H.6 Messiness factors
Messiness factor labels were collected from 2 internal contractors and one researcher.
@@ H.6 Messiness factors
Not purely automatic scoring Does this task NOT have an entirely automatic scoring function? Specifically, for tasks implemented in the task standard: 1. There are no branches in which the scoring function returns None AND 2. The task provides manual scoring instructions in those cases AND 3. There are no clauses about obeying rules that are not actually checked in the automatic scoring. (e.g A family which had the instruction ”you may not attempt to reverse engineer X’” but this is not checked in the automatic scoring function)
@@
H.7 Randomizing file names
@@ H.7.2 Results
H.8 Correlation in performance between agents
In Figure 22, we report the correlation matrix of per-task success rates between each pair of models.
Excess success rates
Excess success rates ($\frac{S_{observed}-S_{predicted}}{S_{predicted}}$) are a metric for how much better (or worse) an AI agent performed compared to what we would expect, given a task’s length and that model’s ability (expected success can be seen in Figure 4).
In Figure 23, we report the correlation matrix for the excess success rate $({S_{observed}-S_{predicted}})/{S_{predicted}}$ (where $S_{predicted}$ is the success rate predicted from the model’s time horizon). While the correlations for excess success rate are lower than the correlation of raw task success rate (0.38 instead of 0.71), the fact that it is positive suggests additional factors that explain model success rates across tasks which are common across models.
Figure 22: Correlation matrix of observed success rates across all models and tasks.
Figure 23: Correlation matrix of excess success rates (defined by $\frac{S_{observed}-S_{predicted}}{S_{predicted}}$) across all models and tasks.
@@ H.8 Correlation in performance between agents
Appendix I Code
Code and data to reproduce some of the core figures in this paper will be provided in the supplementary material; the full codebase is infeasible to publicize and anonymize but the authors can provide code for individual additional figures on reviewer request.
Appendix J Author contributions
Acknowledgments and Disclosure of Funding